OpenStudy (anonymous):

Convert the following degree measures to radians (leaves π in your answer) -370°

OpenStudy (agent0smith):

To convert from degrees to radians, \[\large \theta ^{o} \times \frac{ \pi }{ 180^{o} } \]

OpenStudy (agent0smith):

Since pi is equal to 180 degrees.

OpenStudy (anonymous):

and -370/180=?

OpenStudy (anonymous):

-37pi/18

OpenStudy (agent0smith):

\[\large \frac{ 370 }{180 } = \frac{ 37 }{ 18 }\]

OpenStudy (anonymous):

ohh ,which to click as best response?

OpenStudy (agent0smith):

Pick the one that'll help you solve a similar question in the future.

OpenStudy (anonymous):

tnx Pranav_Gupta and agent0smith